OpenStudy (anonymous):

Find the inverse Laplace transform of F(s)=ln ((2s^2+8)(s+3)^5)/((s^2-16)(s+2)(s+3)^4)

OpenStudy (freckles):

that is ln( that whole thing)?

OpenStudy (freckles):

if so I bet you will need these laws ln(ab)=ln(a)+ln(b) and ln(a/b)=ln(a)-ln(b)

OpenStudy (anonymous):

yes it is

OpenStudy (freckles):

So apply those rules

OpenStudy (freckles):

and then I think the next step would be to differentiate

OpenStudy (anonymous):

\[F(s)=\ln\frac{ (2s^2+8)(s+3)^5 }{ (s^2-16)(s+2)(s+3)^4} \]

OpenStudy (anonymous):

how do you differentiate

OpenStudy (freckles):

example: \[F(s)=\ln(\frac{ s+2}{s-5})=\ln(s+2)-\ln(s-5) \\ \frac{dF}{ds}=\frac{1}{s+2}-\frac{1}{s-5} \\ L(tf(t))=-\frac{dF}{ds}=-(\frac{1}{s+2}-\frac{1}{s-5}) \\ t f(t)=-(e^{-2t}-e^{5t})\]

OpenStudy (freckles):

We first need to expand the ln thing using the rules I mentioned above

OpenStudy (freckles):

\[F(s)=\ln(2s^2+8)+5 \ln(s+3)-\ln(s^2-16)-\ln(s+2)-4 \ln(s+3)\] We could rewrite the part that is the ln(s^2-16) as ln((s-4)(s+4)) as ln(s-4)+ln(s+4)

OpenStudy (freckles):

\[F(s)=\ln(2s^2+8)+5 \ln(s+3)-(\ln(s+4)+\ln(s-4))-\ln(s+2)-4 \ln(s+3) \] Now differentiate F w.r.t s

OpenStudy (anonymous):

\[\frac{ dF }{ ds }=\frac{ 4s }{2s^2+8 }+\frac{ 5 }{ s+3 }-\frac{ 1 }{ s-4 }+\frac{ 1 }{ s-4}-\frac{ 1 }{ s+2 }-\frac{ 4 }{ s+3}\] does that look correct?

OpenStudy (freckles):

almost

OpenStudy (freckles):

just one sign off

OpenStudy (freckles):

there is a -( ln(s+4)+ln(s-4))

OpenStudy (freckles):

-ln(s+4)-ln(s-4)

OpenStudy (freckles):

\[\frac{ dF }{ ds }=\frac{ 4s }{2s^2+8 }+\frac{ 5 }{ s+3 }-\frac{ 1 }{ s-4 }-\frac{ 1 }{ s-4}-\frac{ 1 }{ s+2 }-\frac{ 4 }{ s+3}\]

OpenStudy (freckles):

now we know L(tf(t))=-dF/ds

OpenStudy (freckles):

\[-\frac{dF}{ds}=-(\frac{4}{2} \frac{s}{s^2+4}+5 \frac{1}{s+3}- \frac{1}{s+4}-\frac{1}{s-4}-\frac{1}{s+2}-4 \frac{1}{s+3})\]

OpenStudy (freckles):

and i missed that there was another sign off

OpenStudy (freckles):

so i corrected it on this newer equation

OpenStudy (freckles):

so you are ready to transform now

OpenStudy (anonymous):

so where did u get all that 4/2 and the rest from

OpenStudy (freckles):

I didn't get anything new

OpenStudy (freckles):

\[\frac{4s}{2s^2+8}=\frac{4 \cdot s}{2 \cdot s^2+ 2 \cdot 4} =\frac{4}{2} \frac{s}{s^2+4}\]

OpenStudy (freckles):

I just factored 4 out on top and factored 2 out on bottom for the first fraction

OpenStudy (anonymous):

before we transform, do we need to use partial fractions?

OpenStudy (freckles):

nope not for this there is nothing really to break down

OpenStudy (freckles):

http://www.math.ttu.edu/~gilliam/ttu/s14/chapter_4_lecture_notes.pdf page 27 is helpful

OpenStudy (freckles):

there is only a couple of rows you need from that table to transform this

OpenStudy (anonymous):

\[\frac{ -4 }{ 2 }\cos(2t)+5e^{-3t}+e^{-4t}+e^{4t}+e^{-2t}+4e^{-3t}\] is that the correct transform?

OpenStudy (freckles):

and you also need to transform the left side

OpenStudy (freckles):

and also 4/2 is just 2

OpenStudy (freckles):

remember we had -dF/ds

OpenStudy (freckles):

according to that chart it can be transformed to t*f(t)

OpenStudy (freckles):

\[t \cdot f(t)=-2 \cos(2t)+5e^{-3t}+e^{-4t}+e^{4t}+e^{-2t}+4e^{-3t}\]

OpenStudy (freckles):

now you can solve for f(t)

OpenStudy (anonymous):

\[(-1)^{-2}\cos(2t)\] does the first term by chance look like anything close to that

OpenStudy (freckles):

why that (-1)^(-2) is 1/(-1)^2=1/1=1

OpenStudy (freckles):

first term i see after dividing t on both sides is \[\frac{-2}{t} \cos(2t)\]

OpenStudy (anonymous):

so my professor asked us to use this line from the table to transform\[t^n f(t) \Rightarrow (-1)^nF^{(n)}(s)\] am not sure how to use that

OpenStudy (freckles):

or \[-2t^{-1}\cos(2t)\]

OpenStudy (freckles):

we actually already used that

OpenStudy (freckles):

\[t^{1}f(t) => (-1)^1 F^{(1)}(s)=-\frac{dF}{ds}\]

OpenStudy (freckles):

that is why we found the derivative of F and then multiplied it by -1

OpenStudy (freckles):

so we could use that the transformation of that is t*f(t)

OpenStudy (freckles):

\[t^nf(t)=>(-1)^n \frac{d^nF}{ds^n}\]

OpenStudy (freckles):

\[t^1f(t)=tf(t)=> (-1)^1 \frac{d^1F}{ds^1}=(-1)\frac{dF}{ds} \\ t^2f(t)=> (-1)^2 \frac{d^2F}{ds^2}=\frac{d^2F}{ds^2} \\ t^3 f(t)=> (-1)^3\frac{d^3F}{ds^3}=-\frac{d^3F}{ds^3}\] just trying to show some examples of that thing I started in general

OpenStudy (freckles):

\[F^{(1)}(s) \text{ is the first derivative of } F \text{ w.r.t. } s\]

OpenStudy (freckles):

recall we found that

OpenStudy (freckles):

and multiplied it by -1?

OpenStudy (anonymous):

okay that now makes sense because before we did so many steps together that i missed out on what was going on. now i get it. so basically we just divide through by t

OpenStudy (freckles):

yep since we have tf(t) on the left from the -dF/ds part